OpenStudy (anonymous):

1/infinity = 0 , so infinity *0 =?

OpenStudy (lgbasallote):

\[\infty \times 0\] is an indeterminate form

OpenStudy (anonymous):

there is no such number as \(\frac{1}{\infty}\)

OpenStudy (lgbasallote):

it's supposed to be 1 according to your thingy...but it's indeterminate

OpenStudy (anonymous):

it approaches 0

OpenStudy (anonymous):

It never actually gets there

OpenStudy (anonymous):

while....doing some prob... we used to get...infinity

OpenStudy (anonymous):

mainly...with quatum mechanics....

OpenStudy (anonymous):

similarly, there is no such number as \(0\times \infty\)

OpenStudy (anonymous):

1/infinity = 0 , so infinity *0 =

OpenStudy (lgbasallote):

that depends..usually you have to manipulate it to become 0/0 or ifty/infty then use l'hospital

OpenStudy (lgbasallote):

http://en.wikipedia.org/wiki/Indeterminate_form

OpenStudy (anonymous):

same way there is no number for 2 * infinity

OpenStudy (anonymous):

@009infinity

OpenStudy (anonymous):

0/0

OpenStudy (anonymous):

infinity/infinity

OpenStudy (anonymous):

0 & infinity are two awatar of god if u know about god u can know about that.

OpenStudy (anonymous):

at the risk of repeating myself, \(\infty\) is not a number, and therefore neither is \(0\times \infty\) or \(\frac{1}{\infty}\)

OpenStudy (anonymous):

if any one knows me he knows the answer :-)

OpenStudy (anonymous):

@Yahoo! come on

OpenStudy (anonymous):

thxx